# Unraveling the Age Estimation Puzzle: Comparative Analysis of Deep Learning Approaches for Facial Age Estimation

### v1

@@ 5 Discussion and Conclusions
In this paper, we aimed to establish a fair comparison framework for evaluating various approaches for age estimation. We conducted a comprehensive analysis on seven different datasets, namely AgeDB [20], AFAD [21], CACD2000 [4], CLAP2016 [1], FG-NET [15], MORPH [23], and UTKFace [28], comparing the models based on their Mean Absolute Error (MAE). To determine if any method outperformed the others, we employed the Friedman test and the Nemenyi critical difference test.
When incorporating pre-training on the IMDB-WIKI dataset, we observed no significant improvement in the age estimation results compared to the baseline cross-entropy method. It is widely recognized that as the amount of available data increases, the cross-entropy loss becomes asymptotically optimal for learning the true posterior. Hence, we contend that with the pre-training, we are approaching this asymptotic regime. Previously published results report continuous performance improvements over time (as depicted in Fig. 1), however, our findings challenge these claims. We argue that the reported improvements can be attributed to either the random data splitting strategy or hyperparameter tuning aimed at achieving the best test performance.
When employing random model initialization, we observed some improvement over the baseline cross-entropy on small datasets. Specifically, the Mean-Variance loss [22], OR-CNN [21], and DLDL [12] demonstrated significant improvements. These improvements can be attributed to the regularization provided by these methods.
Furthermore, our analysis of the data preparation pipeline revealed that factors such as the extent of facial coverage, input resolution, and the facial alignment procedure had a more significant impact on the achieved results compared to the choice of the age estimation method itself. Moreover, we suggest that for age estimation tasks outside of the low-data regime, designing specialized methods may not be necessary. The standard approach of utilizing cross-entropy loss is sufficient in such cases. Based on our results, we emphasize the necessity of employing standardized subject-exclusive data splitting when comparing different methods. To facilitate reproducibility and future comparisons, we have made the splits, as well as our implementation, publicly available.
Backbone
@@
AgeDB AFAD CACD2000 CLAP2016 MORPH UTKFace
ResNet-50 5.81 3.14 3.96 4.49 2.81 4.38
EfficientNet-B4 5.76 3.20 4.00 4.06 2.87 4.23
ViT-B-16 9.07 4.04 6.22 8.55 4.35 6.88
VGG-16 6.02 3.22 3.92 4.65 2.88 4.64
Training Dataset Evaluation Dataset AgeDB AFAD CACD2000 CLAP2016 FG-NET MORPH UTKFace Method Init. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. AgeDB Cross-Entropy 5.81 7.20 7.65 7.83 12.61 14.70 5.90 8.10 8.73 6.83 10.86 12.41 10.82 16.27 18.87 4.83 6.74 6.93 8.45 11.88 11.82 Regression 6.23 6.54 7.60 8.05 12.13 14.19 6.76 7.56 8.74 8.32 9.95 12.32 10.56 13.83 18.00 5.64 6.66 6.94 9.42 10.42 12.14 OR-CNN [21] 5.78 6.51 7.52 7.47 11.86 13.80 6.05 7.71 8.55 6.64 10.18 11.74 9.74 13.82 17.40 4.73 6.61 7.05 8.20 10.75 11.41 DLDL [12] 5.80 6.95 7.46 7.81 1.85 14.92 5.99 7.96 8.52 6.51 10.84 11.48 9.23 15.63 16.95 4.74 6.34 6.82 7.97 11.87 11.24 DLDL-v2 [13] 5.80 6.87 7.58 7.61 12.98 16.19 5.91 8.15 8.61 6.48 10.88 12.50 9.91 15.36 19.01 4.92 7.53 7.20 7.97 11.61 12.30 SORD [9] 5.81 6.93 7.58 7.81 12.80 15.42 5.96 7.90 8.77 6.61 10.37 12.22 9.72 14.76 17.85 4.76 6.58 7.14 8.12 11.53 11.60 Mean-Var. [22] 5.85 6.69 7.33 7.26 12.40 14.35 6.00 7.89 8.35 6.70 10.50 11.90 10.55 14.32 17.43 4.99 6.87 7.33 8.25 10.77 11.64 Unimodal [17] 5.90 7.11 15.49 8.37 13.11 20.87 6.22 8.24 16.11 6.73 11.14 21.31 10.15 16.13 32.77 4.84 6.78 17.42 8.23 11.86 23.09 AFAD Cross-Entropy 15.70 17.31 18.05 3.14 3.17 3.32 9.54 11.18 11.21 8.96 10.23 10.32 10.92 11.38 11.96 6.80 6.83 8.19 12.10 13.07 13.29 Regression 13.67 15.91 17.21 3.17 3.16 3.30 8.72 10.51 10.72 8.33 9.91 10.02 11.20 11.89 12.35 6.27 7.34 7.99 11.23 12.83 12.96 OR-CNN [21] 12.08 15.65 16.72 3.16 3.17 3.28 8.87 11.05 10.89 7.85 9.73 9.92 10.63 11.94 12.58 6.68 6.85 7.81 10.50 12.43 12.74 DLDL [12] 14.12 15.70 17.21 3.14 3.16 3.25 9.40 10.70 11.06 8.68 9.54 9.98 11.31 11.64 12.07 7.04 6.75 7.82 11.52 12.43 12.82 DLDL-v2 [13] 13.90 16.33 17.78 3.15 3.17 3.28 9.46 10.68 11.02 8.60 9.76 10.32 10.83 11.81 12.64 6.92 6.79 7.94 11.29 12.61 13.18 SORD [9] 14.30 16.08 17.49 3.14 3.15 3.24 9.45 10.70 11.09 8.64 9.79 10.10 11.21 11.63 12.19 6.87 6.82 7.93 11.59 12.79 13.10 Mean-Var. [22] 12.54 15.07 16.68 3.16 3.16 3.26 8.98 10.33 10.75 7.93 9.33 9.78 10.96 12.24 12.43 6.61 6.76 7.88 10.57 12.00 12.62 Unimodal [17] 13.99 15.89 20.97 3.20 3.24 9.30 9.23 10.68 14.56 8.64 9.79 14.51 11.31 11.83 18.29 7.07 7.32 12.53 11.26 12.33 17.47 CACD2000 Cross-Entropy 9.66 11.84 10.60 10.70 8.50 13.08 3.96 4.59 4.89 8.42 8.64 10.51 17.45 23.64 20.86 7.21 12.20 10.39 11.16 11.38 12.61 Regression 10.91 10.44 10.76 10.23 7.23 11.66 4.06 4.52 4.83 8.84 7.75 9.98 17.55 19.50 19.60 8.61 8.81 11.79 11.34 10.38 11.78 OR-CNN [21] 10.43 11.02 11.85 9.66 9.48 12.17 4.01 4.60 4.74 8.57 8.85 10.29 18.47 24.32 20.85 7.52 10.04 11.05 11.17 12.30 12.27 DLDL [12] 9.84 10.79 11.28 10.09 9.30 13.20 3.96 4.42 4.76 8.39 8.49 9.99 18.38 18.99 21.52 7.27 9.16 11.01 11.19 11.94 12.27 DLDL-v2 [13] 9.90 12.31 11.20 8.03 11.50 11.51 3.96 4.57 4.69 7.67 8.88 9.43 18.11 22.89 19.02 7.20 13.46 9.73 10.52 12.32 11.47 SORD [9] 9.77 10.90 11.04 10.35 9.55 11.95 3.96 4.42 4.70 8.38 8.51 9.89 18.05 20.84 21.73 7.23 8.98 11.59 11.18 12.06 12.22 Mean-Var. [22] 10.81 11.42 10.83 9.71 10.82 11.49 4.07 4.60 4.78 8.88 9.20 10.08 20.48 22.68 20.14 8.14 12.59 11.72 11.74 12.29 12.23 Unimodal [17] 10.46 11.04 46.26 10.63 9.85 25.74 4.10 4.73 37.41 9.19 8.92 30.96 19.37 19.75 15.84 8.94 11.64 32.63 11.89 11.75 32.98 CLAP2016 Cross-Entropy 7.35 10.15 12.26 5.41 7.03 5.34 6.65 8.11 9.11 4.49 5.96 8.73 5.92 9.28 12.02 4.96 6.61 6.90 5.74 7.21 8.58 Regression 7.51 8.52 11.74 6.07 5.19 5.95 6.86 7.24 9.45 4.65 4.77 7.89 4.85 6.31 10.14 5.09 5.49 8.83 6.02 5.93 8.66 OR-CNN [21] 6.83 8.74 11.24 5.83 5.92 5.44 6.73 7.25 8.65 4.13 4.60 7.38 5.09 6.47 9.22 4.92 5.78 6.52 5.43 5.95 7.68 DLDL [12] 7.20 9.33 11.39 5.57 6.90 5.85 6.85 7.64 9.26 4.18 5.10 7.39 5.26 7.44 9.18 4.89 5.92 6.52 5.51 6.37 7.87 DLDL-v2 [13] 7.14 9.42 12.36 5.47 5.95 6.45 6.69 7.99 9.34 4.23 4.87 8.52 5.22 7.04 8.75 4.85 6.04 7.29 5.53 6.12 8.23 SORD [9] 7.19 9.60 12.16 5.47 7.74 6.62 6.63 8.09 9.66 4.27 5.34 7.81 5.59 7.77 7.62 4.92 6.01 6.62 5.48 6.46 8.08 Mean-Var. [22] 7.08 9.16 12.58 5.18 6.30 5.38 6.64 7.37 9.94 4.28 4.87 7.95 5.45 6.69 11.14 4.96 7.38 7.49 5.52 6.16 8.65 Unimodal [17] 7.01 9.77 20.71 5.58 6.10 5.54 6.47 8.20 13.08 4.17 5.39 13.83 5.13 6.39 15.13 4.80 6.05 10.02 5.44 6.67 15.27 MORPH Cross-Entropy 9.66 11.73 12.63 6.69 7.78 10.36 8.53 10.83 10.11 6.90 8.96 10.64 9.45 11.96 15.38 2.81 2.96 3.01 8.97 10.81 11.92 Regression 10.48 12.99 12.56 6.60 6.65 10.66 9.82 11.47 9.68 7.83 9.27 10.67 9.24 10.13 16.69 2.83 2.74 2.97 9.40 10.97 12.06 OR-CNN [21] 9.35 11.65 12.82 6.78 7.78 11.81 8.39 11.34 10.23 6.84 8.73 11.05 9.58 11.09 17.47 2.83 2.85 2.99 8.82 10.37 12.06 DLDL [12] 9.41 12.00 12.66 6.58 7.78 11.76 8.58 11.92 10.10 6.85 9.26 11.15 9.44 11.43 16.94 2.81 2.92 2.98 8.80 10.81 12.46 DLDL-v2 [13] 9.79 11.49 12.68 6.60 8.22 12.45 8.79 10.98 9.81 6.98 8.98 11.22 9.52 11.63 17.57 2.82 2.93 3.00 8.97 10.70 12.47 SORD [9] 9.48 11.84 12.73 6.54 7.91 11.19 8.73 11.18 10.13 6.84 8.99 10.72 9.34 11.08 15.90 2.81 2.91 2.99 8.83 10.85 11.97 Mean-Var. [22] 9.70 11.62 12.93 6.68 7.81 10.41 8.65 10.59 10.11 7.03 8.80 10.56 9.51 11.45 15.81 2.83 2.89 2.95 8.94 10.59 11.95 Unimodal [17] 9.93 12.31 17.44 6.63 7.04 8.18 8.68 10.11 12.03 7.19 8.95 12.38 9.80 12.17 17.83 2.78 2.90 8.66 9.07 10.75 15.45 UTKFace Cross-Entropy 6.61 8.88 9.58 5.51 6.42 6.75 6.56 9.10 8.98 4.82 7.34 7.50 4.78 6.62 7.64 5.09 6.61 7.35 4.38 4.75 5.32 Regression 7.01 7.79 8.83 5.96 6.26 6.43 6.77 7.87 8.61 5.24 5.93 6.67 4.41 5.07 7.27 5.41 5.95 6.71 4.72 4.53 5.34 OR-CNN [21] 6.71 8.29 8.75 5.56 6.74 6.52 6.61 8.89 8.37 4.95 6.79 6.70 4.54 5.71 6.55 5.26 6.07 6.76 4.40 4.43 5.15 DLDL [12] 6.65 8.60 9.00 5.42 6.68 6.19 6.52 9.01 8.84 4.81 7.19 7.46 4.85 5.87 7.28 5.16 6.25 7.03 4.39 4.66 5.30 DLDL-v2 [13] 6.79 8.43 8.91 5.45 6.65 5.82 6.61 9.39 8.50 4.98 7.27 6.85 4.79 5.93 7.44 5.46 6.24 6.79 4.42 4.60 5.19 SORD [9] 6.61 8.96 9.11 5.42 7.18 6.32 6.52 9.42 8.69 4.82 7.87 7.18 4.83 6.15 7.55 5.14 6.36 7.28 4.36 4.68 5.25 Mean-Var. [22] 6.79 8.36 8.53 5.41 6.54 6.32 6.55 8.55 8.32 5.04 6.81 6.32 5.05 6.30 6.90 5.37 6.15 6.39 4.42 4.57 5.05 Unimodal [17] 6.68 8.66 22.42 5.35 7.68 16.64 6.58 9.28 17.17 4.86 7.60 18.83 4.55 6.25 22.98 5.22 5.96 16.44 4.47 4.78 21.01

### v3

@@ 6 Discussion and Conclusions
In this paper, we aimed to establish a fair comparison framework for evaluating various approaches for age estimation. We conducted a comprehensive analysis on seven different datasets, namely AgeDB [21], AFAD [22], CACD2000 [4], CLAP2016 [1], FG-NET [16], MORPH [24], and UTKFace [30], comparing the models based on their Mean Absolute Error (MAE). To determine if any method outperformed the others, we employed the Friedman test and the Nemenyi critical difference test. When pre-training the models on a large dataset, we did not observe any statistically significant improvement by using the specialized loss functions designed for age estimation. With random model initialization, we observed some improvement over the baseline cross-entropy on small datasets. Specifically, for Mean-Variance loss [23], OR-CNN [22], and DLDL [13]. These improvements can be attributed to implicit regularization provided by these methods.
Previously published results reported continuous performance improvements over time (as depicted in Fig. 1). Our findings challenge these claims. We argue that the reported improvements can be attributed to either the random data splitting strategy or hyperparameter tuning to achieve the best test set performance. Our analysis of the data preparation pipeline revealed that factors such as the extent of facial coverage or input resolution exert a more significant impact on the results than the choice of the age estimation specific loss function. Guided by these findings, we use the FaRL [31] model as a backbone for age estimation and demonstrated its effectiveness. In summary:
  • We show that existing evaluation practices in age estimation do not provide a consistent comparison of the state-of-the-art methods. We define a proper evaluation protocol which addresses the issue.
  • We show that improvements in age estimation results over recent years can not be attributed to the specialized loss functions introduced in [22, 13, 14, 10, 23, 18], as is claimed in the published literature.
  • Using the insight gained from analyzing different components of the age estimation pipeline, we construct a prediction model with the FaRL [31] backbone and demonstrate its effectiveness.
  • To facilitate reproducibility and simple future comparisons, we have made our implementation framework and the exact data splits publicly available.
Acknowledgment
This research was supported by the Grant Agency of the Czech Technical University in Prague, grant No. SGS23/176/OHK3/3T/13 and by the Grant agency of the Ministry of Interior Czech Republic, project FACIS grant. No. VJ02010041
Training Dataset Evaluation Dataset AgeDB AFAD CACD2000 CLAP2016 FG-NET MORPH UTKFace Method Init. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. IMDB Imag. Rand. AgeDB Cross-Entropy 5.81 7.20 7.65 7.83 12.61 14.70 5.90 8.10 8.73 6.83 10.86 12.41 10.82 16.27 18.87 4.83 6.74 6.93 8.45 11.88 11.82 Regression 6.23 6.54 7.60 8.05 12.13 14.19 6.76 7.56 8.74 8.32 9.95 12.32 10.56 13.83 18.00 5.64 6.66 6.94 9.42 10.42 12.14 OR-CNN [22] 5.78 6.51 7.52 7.47 11.86 13.80 6.05 7.71 8.55 6.64 10.18 11.74 9.74 13.82 17.40 4.73 6.61 7.05 8.20 10.75 11.41 DLDL [13] 5.80 6.95 7.46 7.81 1.85 14.92 5.99 7.96 8.52 6.51 10.84 11.48 9.23 15.63 16.95 4.74 6.34 6.82 7.97 11.87 11.24 DLDL-v2 [14] 5.80 6.87 7.58 7.61 12.98 16.19 5.91 8.15 8.61 6.48 10.88 12.50 9.91 15.36 19.01 4.92 7.53 7.20 7.97 11.61 12.30 SORD [10] 5.81 6.93 7.58 7.81 12.80 15.42 5.96 7.90 8.77 6.61 10.37 12.22 9.72 14.76 17.85 4.76 6.58 7.14 8.12 11.53 11.60 Mean-Var. [23] 5.85 6.69 7.33 7.26 12.40 14.35 6.00 7.89 8.35 6.70 10.50 11.90 10.55 14.32 17.43 4.99 6.87 7.33 8.25 10.77 11.64 Unimodal [18] 5.90 7.11 15.49 8.37 13.11 20.87 6.22 8.24 16.11 6.73 11.14 21.31 10.15 16.13 32.77 4.84 6.78 17.42 8.23 11.86 23.09 FaRL + MLP 5.64 - - 7.82 - - 7.41 - - 9.32 - - 9.15 - - 4.73 - - 9.92 - - AFAD Cross-Entropy 15.70 17.31 18.05 3.14 3.17 3.32 9.54 11.18 11.21 8.96 10.23 10.32 10.92 11.38 11.96 6.80 6.83 8.19 12.10 13.07 13.29 Regression 13.67 15.91 17.21 3.17 3.16 3.30 8.72 10.51 10.72 8.33 9.91 10.02 11.20 11.89 12.35 6.27 7.34 7.99 11.23 12.83 12.96 OR-CNN [22] 12.08 15.65 16.72 3.16 3.17 3.28 8.87 11.05 10.89 7.85 9.73 9.92 10.63 11.94 12.58 6.68 6.85 7.81 10.50 12.43 12.74 DLDL [13] 14.12 15.70 17.21 3.14 3.16 3.25 9.40 10.70 11.06 8.68 9.54 9.98 11.31 11.64 12.07 7.04 6.75 7.82 11.52 12.43 12.82 DLDL-v2 [14] 13.90 16.33 17.78 3.15 3.17 3.28 9.46 10.68 11.02 8.60 9.76 10.32 10.83 11.81 12.64 6.92 6.79 7.94 11.29 12.61 13.18 SORD [10] 14.30 16.08 17.49 3.14 3.15 3.24 9.45 10.70 11.09 8.64 9.79 10.10 11.21 11.63 12.19 6.87 6.82 7.93 11.59 12.79 13.10 Mean-Var. [23] 12.54 15.07 16.68 3.16 3.16 3.26 8.98 10.33 10.75 7.93 9.33 9.78 10.96 12.24 12.43 6.61 6.76 7.88 10.57 12.00 12.62 Unimodal [18] 13.99 15.89 20.97 3.20 3.24 9.30 9.23 10.68 14.56 8.64 9.79 14.51 11.31 11.83 18.29 7.07 7.32 12.53 11.26 12.33 17.47 FaRL + MLP 16.41 - - 3.12 - - 10.95 - - 8.57 - - 12.24 - - 6.62 - - 11.64 - - CACD2000 Cross-Entropy 9.66 11.84 10.60 10.70 8.50 13.08 3.96 4.59 4.89 8.42 8.64 10.51 17.45 23.64 20.86 7.21 12.20 10.39 11.16 11.38 12.61 Regression 10.91 10.44 10.76 10.23 7.23 11.66 4.06 4.52 4.83 8.84 7.75 9.98 17.55 19.50 19.60 8.61 8.81 11.79 11.34 10.38 11.78 OR-CNN [22] 10.43 11.02 11.85 9.66 9.48 12.17 4.01 4.60 4.74 8.57 8.85 10.29 18.47 24.32 20.85 7.52 10.04 11.05 11.17 12.30 12.27 DLDL [13] 9.84 10.79 11.28 10.09 9.30 13.20 3.96 4.42 4.76 8.39 8.49 9.99 18.38 18.99 21.52 7.27 9.16 11.01 11.19 11.94 12.27 DLDL-v2 [14] 9.90 12.31 11.20 8.03 11.50 11.51 3.96 4.57 4.69 7.67 8.88 9.43 18.11 22.89 19.02 7.20 13.46 9.73 10.52 12.32 11.47 SORD [10] 9.77 10.90 11.04 10.35 9.55 11.95 3.96 4.42 4.70 8.38 8.51 9.89 18.05 20.84 21.73 7.23 8.98 11.59 11.18 12.06 12.22 Mean-Var. [23] 10.81 11.42 10.83 9.71 10.82 11.49 4.07 4.60 4.78 8.88 9.20 10.08 20.48 22.68 20.14 8.14 12.59 11.72 11.74 12.29 12.23 Unimodal [18] 10.46 11.04 46.26 10.63 9.85 25.74 4.10 4.73 37.41 9.19 8.92 30.96 19.37 19.75 15.84 8.94 11.64 32.63 11.89 11.75 32.98 FaRL + MLP 11.32 - - 9.08 - - 3.96 - - 8.57 - - 19.63 - - 6.56 - - 11.27 - - CLAP2016 Cross-Entropy 7.35 10.15 12.26 5.41 7.03 5.34 6.65 8.11 9.11 4.49 5.96 8.73 5.92 9.28 12.02 4.96 6.61 6.90 5.74 7.21 8.58 Regression 7.51 8.52 11.74 6.07 5.19 5.95 6.86 7.24 9.45 4.65 4.77 7.89 4.85 6.31 10.14 5.09 5.49 8.83 6.02 5.93 8.66 OR-CNN [22] 6.83 8.74 11.24 5.83 5.92 5.44 6.73 7.25 8.65 4.13 4.60 7.38 5.09 6.47 9.22 4.92 5.78 6.52 5.43 5.95 7.68 DLDL [13] 7.20 9.33 11.39 5.57 6.90 5.85 6.85 7.64 9.26 4.18 5.10 7.39 5.26 7.44 9.18 4.89 5.92 6.52 5.51 6.37 7.87 DLDL-v2 [14] 7.14 9.42 12.36 5.47 5.95 6.45 6.69 7.99 9.34 4.23 4.87 8.52 5.22 7.04 8.75 4.85 6.04 7.29 5.53 6.12 8.23 SORD [10] 7.19 9.60 12.16 5.47 7.74 6.62 6.63 8.09 9.66 4.27 5.34 7.81 5.59 7.77 7.62 4.92 6.01 6.62 5.48 6.46 8.08 Mean-Var. [23] 7.08 9.16 12.58 5.18 6.30 5.38 6.64 7.37 9.94 4.28 4.87 7.95 5.45 6.69 11.14 4.96 7.38 7.49 5.52 6.16 8.65 Unimodal [18] 7.01 9.77 20.71 5.58 6.10 5.54 6.47 8.20 13.08 4.17 5.39 13.83 5.13 6.39 15.13 4.80 6.05 10.02 5.44 6.67 15.27 FaRL + MLP 7.50 - - 4.34 - - 6.57 - - 3.38 - - 4.95 - - 4.47 - - 4.85 - - MORPH Cross-Entropy 9.66 11.73 12.63 6.69 7.78 10.36 8.53 10.83 10.11 6.90 8.96 10.64 9.45 11.96 15.38 2.81 2.96 3.01 8.97 10.81 11.92 Regression 10.48 12.99 12.56 6.60 6.65 10.66 9.82 11.47 9.68 7.83 9.27 10.67 9.24 10.13 16.69 2.83 2.74 2.97 9.40 10.97 12.06 OR-CNN [22] 9.35 11.65 12.82 6.78 7.78 11.81 8.39 11.34 10.23 6.84 8.73 11.05 9.58 11.09 17.47 2.83 2.85 2.99 8.82 10.37 12.06 DLDL [13] 9.41 12.00 12.66 6.58 7.78 11.76 8.58 11.92 10.10 6.85 9.26 11.15 9.44 11.43 16.94 2.81 2.92 2.98 8.80 10.81 12.46 DLDL-v2 [14] 9.79 11.49 12.68 6.60 8.22 12.45 8.79 10.98 9.81 6.98 8.98 11.22 9.52 11.63 17.57 2.82 2.93 3.00 8.97 10.70 12.47 SORD [10] 9.48 11.84 12.73 6.54 7.91 11.19 8.73 11.18 10.13 6.84 8.99 10.72 9.34 11.08 15.90 2.81 2.91 2.99 8.83 10.85 11.97 Mean-Var. [23] 9.70 11.62 12.93 6.68 7.81 10.41 8.65 10.59 10.11 7.03 8.80 10.56 9.51 11.45 15.81 2.83 2.89 2.95 8.94 10.59 11.95 Unimodal [18] 9.93 12.31 17.44 6.63 7.04 8.18 8.68 10.11 12.03 7.19 8.95 12.38 9.80 12.17 17.83 2.78 2.90 8.66 9.07 10.75 15.45 FaRL + MLP 8.40 - - 4.67 - - 7.45 - - 6.21 - - 9.28 - - 3.04 - - 8.93 - - UTKFace Cross-Entropy 6.61 8.88 9.58 5.51 6.42 6.75 6.56 9.10 8.98 4.82 7.34 7.50 4.78 6.62 7.64 5.09 6.61 7.35 4.38 4.75 5.32 Regression 7.01 7.79 8.83 5.96 6.26 6.43 6.77 7.87 8.61 5.24 5.93 6.67 4.41 5.07 7.27 5.41 5.95 6.71 4.72 4.53 5.34 OR-CNN [22] 6.71 8.29 8.75 5.56 6.74 6.52 6.61 8.89 8.37 4.95 6.79 6.70 4.54 5.71 6.55 5.26 6.07 6.76 4.40 4.43 5.15 DLDL [13] 6.65 8.60 9.00 5.42 6.68 6.19 6.52 9.01 8.84 4.81 7.19 7.46 4.85 5.87 7.28 5.16 6.25 7.03 4.39 4.66 5.30 DLDL-v2 [14] 6.79 8.43 8.91 5.45 6.65 5.82 6.61 9.39 8.50 4.98 7.27 6.85 4.79 5.93 7.44 5.46 6.24 6.79 4.42 4.60 5.19 SORD [10] 6.61 8.96 9.11 5.42 7.18 6.32 6.52 9.42 8.69 4.82 7.87 7.18 4.83 6.15 7.55 5.14 6.36 7.28 4.36 4.68 5.25 Mean-Var. [23] 6.79 8.36 8.53 5.41 6.54 6.32 6.55 8.55 8.32 5.04 6.81 6.32 5.05 6.30 6.90 5.37 6.15 6.39 4.42 4.57 5.05 Unimodal [18] 6.68 8.66 22.42 5.35 7.68 16.64 6.58 9.28 17.17 4.86 7.60 18.83 4.55 6.25 22.98 5.22 5.96 16.44 4.47 4.78 21.01 FaRL + MLP 7.16 - - 4.69 - - 6.93 - - 4.02 - - 5.07 - - 4.76 - - 3.87 - -
Supplementary Material
Appendix A Literature Survey
To thoroughly survey the CVPR and ICCV, we adhered to the following procedure. We systematically searched for all papers presented at the CVPR or ICCV conferences from 2013 onwards that encompassed keywords such as "age", "aging", "face", "facial", and "ordinal" in their titles. We excluded papers that primarily focused on face detection, recognition, or editing, as our specific focus was on age estimation. However, we retained papers that addressed the learning of facial representations (such as unsupervised pre-training and clustering), as they could evaluate the quality of their representations on the age estimation task.
Subsequently, we meticulously reviewed the remaining papers to determine whether they centered around age estimation. We documented the datasets employed by these papers and their data partitioning strategies. Additionally, we extended our survey to include age estimation literature from other conferences that were referenced by the aforementioned papers.
Our estimates, based on the surveyed literature, indicate that approximately $70\%$ of the papers that attempt to improve age estimation follow this approach. Only a minority of these papers adequately ablate the impact of the proposed modifications. Most of the remaining papers suggest modifications to the training procedure, backbone, or aspects of the data pipeline.
Appendix B Compared Methods
Classification
Extended Binary Classification
Fixed Distribution Learning
Gao et al. [13] (DLDL) approach the task as multi-class classification. However, they encode the label distribution as a normal distribution centered at the true label. Díaz and Marathe [10] (SORD) approach the task similarly, but encode the label distribution as a double exponential distribution centered at the true label. In a follow up to their work [13], Gao et al. [14] (DLDL-v2) propose to also minimize the difference between {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
the true label $y\in\mathcal{Y}$
the expectation $\mathbb{E}_{\hat{y}\sim f(x)}[\hat{y}]$ of the model output distribution $f(x)$.
Adaptive Distribution Learning
An approach emerging in recent years is not to model a specific distribution, such as normal or double exponential distribution, but instead, to constrain the model by some statistical measure or a condition. Pan et al. [23] (Mean-Variance) approach the task as standard multi-class classification, but design a loss function that {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
minimizes the squared difference between the expectation $\mathbb{E}_{\hat{y}\sim f(x)}[\hat{y}]$ and the true label $y\in\mathcal{Y}$
minimizes the variance $\mathbb{E}_{\bar{y}\sim f(x)}\left[(\bar{y}-\mathbb{E}_{\hat{y}\sim f(x)}\left[\hat{y}\right])^{2}\right]$ of the model output distribution $f(x)$. Similarly, Li et al. [18] (Unimodal) design a loss function {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
which constrains the model to output unimodal distributions
concentrates the output distribution around the true label $y\in\mathcal{Y}$.
Note on Prediction Strategy
Note that for all methods which model the posterior distribution $p(y|x)$, namely {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
cross-entropy
DLDL [13]
DLDL-v2 [14]
SORD [10]
Mean-Variance loss [23]
Unimodal loss [18] , we use the optimal plugin Bayes predictor for MAE loss, i.e., we predict $\arg\min_{y}\mathbb{E}_{\hat{y}\sim f(x)}[|y-\hat{y}|]$. For regression, we use the absolute error as the loss function.
Appendix C Additional Comments
Pre-training
For some experiments, we pre-train the models on IMDB-WIKI. However, it is important to note that the labels (bounding box, identity, age) in the IMDB-WIKI dataset are known to be noisy. To mitigate this problem, Lin et al. [20], and Franc and Čech [12] attempted to clean the labels. To assess the quality of these labels, we trained ResNet-50 models on the dataset using the labels proposed by Lin et al. [20] and Franc and Čech [12], and evaluated the models’ performance on the other datasets [21, 1, 4, 24, 30, 22]. The results are presented in Tab. 7. Both models achieved similar results, so the choice of labels between [20, 12] is in our opinion arbitrary. Due to a slightly lower overall Mean Absolute Error (MAE), we decided to use the labels from Franc and Čech [12] for model pre-training in this paper.
Evaluation
@@ Appendix C Additional Comments
Annotations
EM-CNN [12] FP-AGE [20]
AgeDB 6.44 6.30
AFAD 6.86 7.23
CACD2000 5.81 5.90
CLAP2016 6.24 5.53
FG-NET 10.32 6.09
MORPH 4.94 5.30
UTKFace 8.31 6.26
Overall 6.28 6.36
IMDB 4.90 5.15
Comment on Task Uncertainty
The irreducible Bayes error of the age estimation task is contingent on the specific formulation. When estimating the real age from an observation $x_{i}$, the label $y_{i}$ is a realization of the distribution $p(y\,|\,x)$, a random variable. The observation $x_{i}$ does not contain all the necessary information about the person’s genetics, lifestyle, etc., and the Bayes error is non-zero. Interestingly, when estimating the apparent age, the Bayes error can be 0. Specifically, when the label is defined as $y_{i}=\mathbb{E}_{\hat{y}_{i}\sim p(y\,|\,x)}[{\hat{y}_{i}}]$, which is not a random variable. Apparent age is by definition the expected annotation provided by observers. Therefore the observation $x_{i}$ must necessarily contain all the information used to generate the label. The process of dataset collection provides further insight. For AgeDB, AFAD, and CACD2000, the age label is defined as the year when the photo was taken minus the year of birth. Two individuals born on December 31st and January 1st would thus have different age labels, even though they were born just a day apart. The year taken and the year born can also be noisy. For instance, AFAD collects data from a social network with a minimum age requirement of 15, without verifying the actual age of the users. CLAP2016, FG-NET, and MORPH use the legal age as the age label. Still, uncertainty arises from the discrete nature of the label. For UTKFace, the labels are estimated by DEX [25], and manually verified. Therefore, DEX should attain zero error on UTKFace.